Please help me sovle this problem

Question 5. Thejoint probability distribution of two discrete random vari- ables X and Y is given in the following table. Take e to be fixed between -1/4 and 1/4: 1/4-? 1/4+? 1/4+1 1/4-e px(a) 1/2 1/2 py(b) 1/2 1/2 (a) Take e 1/8 and compute cov(X, Y) (b) Take-18 and compute p(X, Y). (c) For which values of e is ?(X, Y) equal to-1, 0,1?
Answer #1

a) taking \epsilon =1/8 ; below is joint probability distribution:

y
x 0 1 Total
0    1/8    3/8    1/2
1    3/8    1/8    1/2
Total    1/2    1/2 1       

marginal distribution of X:

x P(x) xP(x) x^2P(x)
0 0.5000 0.0000 0.0000
1 0.5000 0.5000 0.5000
total 1 0.5 0.5
E(x) = 0.5000
E(x^2) = 0.5000
Var(x) E(x^2)-(E(x))^2 0.2500

marginal distirbution of Y:

y P(y) yP(y) y^2P(y)
0 0.5000 0.0000 0.0000
1 0.5000 0.5000 0.5000
total 1.0000 0.5000 0.5000
E(y) = 0.5000
E(y^2) = 0.5000
Var(y) E(y^2)-(E(y))^2 0.2500

E(XY) = \sum xyP(x,y) =(1/8)*1*1 =1/8 =0.125

hence Cov(X,Y) =E(XY)-E(X)*E(Y) =0.125-0.5*0.5 =-0.125

b)

correlation coefficient =Cov(X,Y)/(Var(X)*Var(Y))1/2 =-0.5

c)

for \rho (X,Y) =-1 ; xshould increase when Y decrease or vice versa ; therefore \epsilon =1/4

\rho(X,Y) =0 ; x and Y should be independent or P(X=1;Y=1) =P(X=1)*P(Y=1) ; solving whcih \epsilon =0

for \rho (X,Y) =1 ; xshould increase when Y increase or vice versa ; therefore \epsilon =-1/4